A 1200 kg car is traveling at 27 m/s around a level curve of radius 110m. What centripetal force must be applied to prevent the car from slipping, and what is the minimum value for the coefficient of friction required to provide the force?
Answer #1

F = m*v^2 / r = 1200 * (27^2)/110 = 7952.727272727272

Friction = u(fr.cofficient ) * N = 7952.727272
N = mg = 1200*9.8
u (fr. cofficient )= 7952.727272/(1200*9.8) = 0.6765
Answer #2

a)

F = m*v^2/r

= 1200*27^2/110 = 7952.73 N

b)

N = m*g = 1200*9.8 = 11760 N

m*v^2/r = u*N

7952.73 = u*11760

u = 0.676
Answer #3

centripetAL FORCE = mv2 / r

                          =795207272 newton

now ,

coefficient of friction = centripetal force/ mg

                               = 0.67625